OpenStudy (anonymous):

Kim walked to the grocery store and back home.The grocery store is 2 5/8 miles from her house. The next day kim walked twice as far to the post office and then back home.How far did she walk in all?

OpenStudy (tkhunny):

Just add them all up. 2(2 5/8 miles) = 4 5/4 miles = 5 1/4 miles -- Grocery Store and Back 2(5 1/4 miles) = 10 2/4 miles -- Post Office and Back (Deliberately failing to simplify the fractional portion.) Total: 5 1/4 miles + 10 2/4 miles = 15 3/4 miles Do another one so I can see you wade through it.